Definition of theoretical probability?
The term "theoretical probability" is used in contrast to the term "experimental probability" to describe what the result of some trial or event should be based on math, versus what it actually is, based on running a simulation or actually performing the task. For example, the theoretical probability that a single standard coin flip results in heads is 1/2. The experimental probability in a single flip would be 1 if it returned heads, or 0 if it returned tails, since the experimental probability only counts what actually happened.
Each of the dice has 6 sides.
There are 6 x 6 = 36 possible combinations.
Of those, the following equal 5:
1 and 4
2 and 3
3 and 2
4 and 1
So the odds of rolling a 5 is
4 chances out of 36
4 / 36 = 1 / 9 That converts to a probability of 0.1111111...
1/26
9/2
Why use standard deviation in psychology?
Measurements of the variable of interest will differ between subjects. As a result the outcome of an experiment may be different from what was expected. The psychologist will want to know whether that difference is purely the result of random variation between subjects or if it is due to some other factor. The standard deviation is a measure for determining how large a difference might be purely random.
Fraction: 4/13
Percentage: About 30%
Steps:
1. Find out the denominator.
There are 52 cards in a standard deck, so the denominator is 52.
2. Find out how many hearts and 7's there are in a deck. (this is the numerator)
There are 13 hearts, and four 7's in a deck. Since one of the card's is the heart of 7, it's been counted twice, so there are 16 cards that are heart's and 7's.
3. Reduce the fraction
It can be reduced to 4/13
4. Turn it into a percentage, if you need to.
Do 4/13, multiply it by 100, and then write it with the "%" sign.
The Binomial distribution has the probability function P(x=r)= nCr p^r (1-p)^(n-r) r=0,1,2,.....,n where nCr = n! / r! (n-r)! n=number of tosses =10 p=probability of a head in a single toss=1/2 r=8,9,10 We have to compute each probability and add them all up. P(r=8) = 10C8 (1/2)^8 (1/2)^2 P(r=9) = 10C9 (1/2)^9 (1/2) P(r=10) = 10C10 (1/2)^10 http://www.pindling.org/Math/Statistics/Textbook/Functions/Binomial/binomial_2_8.htm The above link computes each probability. =0.0439+0.0098+0.0010 = 0.0547 is the required probability. One may also use a calculator to compute these expressions.
How do you get a Probability using two spinners?
i think the probability of getting a vowl of th spinner is:1,2,3,4
What number is the next logical number in the following sequence of numbers 1 1 2 3 5 8 13?
21. This is the Fibonacci series.
the answer will be nothing
3/6 * 3/5 = 6/30 or 1/5 so you have a 20% chance of pulling a white and then black marble.
If there are 4 choices and you randomly choose one, you have a 25% chance that it will be correct, with 75% that it will be wrong. However, if there is one answer that you know is not correct, you can eliminate that one. Then if you choose from the remaining three, you will have increased your chances of getting it right to 33%. That doesn't sound like a whole lot, but it really would help you.
Probability that a girl is chosen = 23/45 = .511 So, the probability that a boy is chosen = 1 - .511 = .489
Firstly, the probability when tossing a coin and getting a head or tail is 1/2, then rolling a die, there are 6 sides so the chance of rolling any number is 1/6, there are 2 chances of rolling greater than 4 ie 5 and 6, so the probability of rolling a 5 or 6 in 1/3, as these are independent events you multiply the probability getting a heads of tails, (1/2) by the probability of rolling a five or six, (1/3) which gives you 1/6 or 0.1666 recurring.
Does a negative z score yield a negative probability value?
no, z score can be negative but a probability is a always positive between 0 and 1.
What is probability of NOT winning if the chances of winning is 0.008 and you try 110 times?
The chance of not winning on one attempt is 1 - 0.008 = 0.992 The chance of not winning over 110 attempts is therefore 0.992110, since you must fail every time. 0.992110 = 0.413 (to three decimal places). (It follows that 0.587 is the chance of winning on at least one attempt.)
1/2, or 50% since you are only asking what the probability of the last outcome is.
What is the probability of throwing a natural number on a dice?
The probability is 1. (aka 100% aka 6/6)